OpenStudy (chris215):

Find the inverse of the function. f(x)=3 cube rt (x/7) -9

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=3\sqrt[3]{\frac{x}{7}}-9 }\) like this?

OpenStudy (chris215):

\[f(x)= \sqrt[3]{x/7}-9\]

OpenStudy (solomonzelman):

oh, so then f(x)=cube rt (x/7) - 9

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=\sqrt[3]{\frac{x}{7}}-9 }\)

OpenStudy (chris215):

yeah

OpenStudy (solomonzelman):

Steps: `1. For convenience replace f(x) with y` (y and f(x) are same thing, just for further manipulations) `2. Switch x and y.` (put x instead of y, and put y instead of x) `3. In new equation, isolate the y.`

OpenStudy (solomonzelman):

(you will then re-write \({\rm f}^{-1}(x)\) instead of y, after step 3, to denote that it is an inverse function)

OpenStudy (solomonzelman):

ok, can you do step 1 for me please? here is the code I am using, if you want `\(\large\color{black}{ \displaystyle f(x)=3\sqrt[3]{\frac{x}{7}}-9 }\)`

OpenStudy (chris215):

\[x=\sqrt[3]{y/7} -9\]

OpenStudy (solomonzelman):

very good, now isolate the y.

OpenStudy (chris215):

idk y but your code wasn't working

OpenStudy (solomonzelman):

it should be, if you just put in text (not equation editor) `\(\large\color{black}{ \displaystyle f(x)=\sqrt[3]{\frac{x}{7}}-9 }\)` gives me \(\large\color{black}{ \displaystyle f(x)=\sqrt[3]{\frac{x}{7}}-9 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x=\sqrt[3]{\frac{y}{7}}-9 }\)

OpenStudy (solomonzelman):

obtained by: `\(\large\color{black}{ \displaystyle x=\sqrt[3]{\frac{y}{7}}-9 }\)`

OpenStudy (solomonzelman):

just copy paste, and put in your a reply. You can modify the code too.

OpenStudy (solomonzelman):

I mean it is fine the way you are doing with equation editor too.

OpenStudy (chris215):

y=7x^3+189x^2+1701x+5103 that's what I got when I solved for y

OpenStudy (chris215):

but it's not one of my answer choices

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x=\sqrt[3]{\frac{y}{7}}-9 }\) steps to isolate the y: \(\rm 1.~~~~Add~~9~~to~~both~~sides.\) \(\rm 2.~~~~Raise~~both~~sides~~to~~the~~3rd~~power.\) your answer is correct

OpenStudy (solomonzelman):

can I see the options, though?

OpenStudy (solomonzelman):

yes, you didn't need to expand;)

OpenStudy (solomonzelman):

leave it as it is, after you raised both sides to the 3rd power, multiply times 7 on both sides, and you are done.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x+9=\sqrt[3]{x/7} }\) \(\large\color{black}{ \displaystyle (x+9)^3=\left(\sqrt[3]{x/7}~\right)^3 }\) \(\large\color{black}{ \displaystyle (x+9)^{3}=\left(\sqrt[\cancel{3}]{x/7}~\right)^{\cancel{3}} }\) \(\large\color{black}{ \displaystyle (x+9)^3=x/7 }\) and then final step to multiply both sides times 7.

OpenStudy (solomonzelman):

hope this is making sense:)

OpenStudy (chris215):

oh ok thanks so much so it would be that last one

OpenStudy (solomonzelman):

yup, that is option D. :D

OpenStudy (solomonzelman):

which simplifies to D^2

OpenStudy (solomonzelman):

jk